OpenStudy (anonymous):

let f(x)=3x-4 and g(x)=-x^2 . evaluate the composite function. (f.f)(2)

OpenStudy (anonymous):

if f.f is f(f(x)), then f.f(2) = 2

OpenStudy (lalaly):

my bad its f(f(x)=3[3x-4]-4=9x-16 f(f(2)=18-16=2